Combined, there are 192 Asians, Africans, Europeans, and Americans in a village. The number of Asians exceeds the number of Africans and Europeans by 59. The difference between the number of Europeans and Americans is 23. If the number of Africans is doubled, their population exceeds the number of Europeans and Americans by 29. Determine the number of Asians, Africans, Europeans, and Americans in this village.
Found 2 solutions by josgarithmetic, ikleyn:
Answer by josgarithmetic(39617)
Answer by ikleyn(52781)

.
Combined, there are 192 Asians, Africans, Europeans, and Americans in a village.
The number of Asians exceeds the number of Africans and Europeans by 59.
The difference between the number of Europeans and Americans is 23.
If the number of Africans is doubled, their population exceeds the number
of Europeans and Americans by 29. Determine the number of Asians, Africans,
Europeans, and Americans in this village.
~~~~~~~~~~~~~

Let x = Asians
    y = Africans
    z = Europeans
    w = Americans


Write equations as you read the problem

    x + y + z + w = 192
    x - y - z     =  59
            z - w =  23
       2y - z - w =  29


Solve using any method you want/(you like).
I used free of charge online solver www.reshish.com.


The solution set is

x = 122
y =  33
z =  30
w =   7

Solved.